OpenStudy (buttermequeasy):

If f(x) = 1 over 3x − 2 and g(x) = −2x − 1, find (f • g)(9) 19 −18 0 −19

OpenStudy (buttermequeasy):

I got -17. ._.

OpenStudy (anonymous):

find \[f(9)\] then \[g(9)\] then multiply

OpenStudy (anonymous):

\[f(9)=\frac{1}{3\times 9-2}=\frac{1}{25}\]

OpenStudy (anonymous):

\[g(9)=-2\times 9-1=-19\] so i don't get any of your choices. is this \[f\times g(9)\] or \[f\circ g(9)\]

OpenStudy (buttermequeasy):

?

OpenStudy (anonymous):

is \[f(x)=\frac{1}{3x-2}\]?

jhonyy9 (jhonyy9):

fxg=(-2x-1)/(3x-2) fxg(9)=(-2*9-1)/(3*9-2)=(-18-1)/(27-2)=-19/25